OpenStudy (anonymous):

Could someone please tell me how to solve this? If h(x) = (f o g) and h(x) =4(x+1)^2, find f(x) and g(x)

OpenStudy (anonymous):

This isn't a very good question, because there is an infinite number of things that f and g could be.

OpenStudy (anonymous):

One possibility is f(x) = 4x^2 and g(x) = x+1 (plug them in to verify) Another possibility is that f(x) = 4x and g(x) = (x+1)^2